Osborn Manufacturing uses a predetermined overhead rate of $19.50 per direct labor-hour. This predetermined rate was based on a cost formula that estimates $259,350 of total manufacturing overhead for an estimated activity level of 13,300 direct labor-hours.

The company incurred actual total manufacturing overhead costs of $253,000 and 12,800 total direct labor-hours during the period.

Required:

1. Determine the amount of underapplied or overapplied manufacturing overhead for the period.

Manufacturing overhead overapplied/underapplied (choose one) by ________.

2. Assuming that the entire amount of the underapplied or overapplied overhead is closed out to cost of goods sold, what would be the effect of the underapplied or overapplied overhead on the company's gross margin for the period?

The gross margin would decrease/increase (choose one) by ________.
Answer #1
Concepts and reason

Job Costing: This type of costing is used when goods are produced based upon the specifications expected from the customer at different situations.

When goods are produced based upon the expectations of the customer is known as Job order.

It is essential for determining the cost of the job. In this case, Job costing is used to estimate the cost of the manufacturing the products as required by the customers. Therefore it is clear that cost for each job differ from one another.

Job costing usually involves in keeping a track of direct costs and indirect costs that are related with the particular job. In this method, overheads that are related to the jobs are allocated based on Direct Labor Hours.

Manufacturing Costs: The costs which do not relate directly with the manufacturing of products are referred to as manufacturing overhead costs or indirect costs.

Fundamentals

Over or under applied overhead:

When the applied overhead cost is more than the actual overhead cost for a particular period, then the overhead is over applied.

When the applied overhead cost is less than the actual overhead cost, then the overhead cost is under applied.

Predetermined overhead rate: The predetermined overhead allocation rate is used to allocate the indirect overhead costs to the units produced.

Cost of Goods sold: Cost of goods sold is the amount that the merchandiser pays for the sales that have been made during the period. Cost of goods sold is otherwise called as cost of sales.

(1)

Calculate applied overhead:

Appliedoverhead=Actualdirectlaborhours×Predeterminedoverheadrate=12,800×$19.5=$249,600\begin{array}{c}\\{\rm{Applied overhead = Actual direct labor hours }} \times {\rm{ Predetermined overhead rate}}\\\\{\rm{ = 12,800 }} \times {\rm{ \$ 19}}{\rm{.5}}\\\\{\rm{ = \$ 249,600}}\\\end{array}

Hence, Applied overhead is $249,600.

Calculate over or under applied overhead:

Underappliedoverhead=AppliedoverheadActualoverhead=$249,600$253,000=$3,400\begin{array}{c}\\{\rm{Under applied overhead = Applied overhead }} - \,{\rm{Actual overhead}}\\\\{\rm{ = \$ 249,600 }} - \,\$ 253,000\\\\ = \,\$ 3,400\\\end{array}

Therefore, under applied manufacturing overhead is $3,400.

(2)

The effect of under applied overhead of $3,400 would be boost the cost of goods sold and reduced the gross margin for O’s Manufacturing company.

Therefore, Gross margin of O’s Manufacturing Company would be decrease by $3,400.

Ans: Part 1

Under applied overhead is $3,400.

Part 2

Gross margin would decrease by $3,400.